what is the relation between focal lenght and distance between screen and lenght in normal and myopic eye


1
Expert's answer
2021-01-24T21:46:49-0500

The relation between focal lenght and distance between screen and image distance of eye can be written as follows:


"\dfrac{1}{f}=\dfrac{1}{u}+\dfrac{1}{v},"

here, "f" is the focal length of the eye, "u" is the distance from the screen to the eye, "v" is the distance from the eye to the image that is focused on retina (since the diameter of eye is 2.4 cm we assume "v=2.4\ cm").

Let's assume that the distance from the eye to the screen is 25 cm (near point of a human eye), than the focal distance for normal eye calculated from the formula above equals 2.19 cm. For the myopic eye the focal distance will be too short, much less then 2.19 cm, so the correction will needed for such eyes.